# Learning in Structured Stackelberg Games

### v1

@@ 3 Online Learning in Structured Stackelberg Games
Consider an online Stackelberg game in which the leader is faced by a sequence of followers. In each round $t\in[T]$, a context $\mathbf{z}_{t}\in\pazocal{Z}$ is revealed to both players. A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best responds to it and the context according to Equation 1. As before, we assume that the leader knows the utilities of all follower types, but the type of follower at time $t$ is not revealed to the leader until after the round is over. In this section, we make no assumptions about either the contexts or the follower types that appear at each timestep. In fact, we handle the case, where they both could be chosen jointly by an adversary.
3.1 Leader’s task in the realizable setting
Consider a set $\pazocal{H}$ (also called a hypothesis class) of candidate functions mapping contexts to follower types, which the leader uses to learn. More formally, for each $h\in\pazocal{H}$, we let $h:\pazocal{Z}\rightarrow\mathscr{K}$. We first consider the realizable setting, where there exists a true mapping $h^{*}\in\pazocal{H}$ such that $h^{*}(\mathbf{z}_{t})=f_{t}$ for all possible $(\mathbf{z}_{t},f_{t})$ pairs. We relax this assumption in Section 3.2. We use $h^{*}$ as a benchmark to measure the leader’s performance over time, and we quantify this via the notion of contextual Stackelberg regret.
@@ 3.1 Leader’s task in the realizable setting
Given a sequence of followers $f_{1},\ldots,f_{T}$ and contexts $\mathbf{z}_{1},\dots\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret is given by:
$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})).$
We let $\pi_{h^{*}}(\mathbf{z})=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ and we assume that there is a fixed and arbitrary tie-breaking rule over the leader strategies. We let $\mathbf{x}_{t}$ for $t\in[T]$ be the mixed strategy played by the leader at time $t$.
The goal of the leader is to learn a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ that minimizes her regret, that is, it achieves utility comparable to the utility she would get if she could (1) optimally predict the follower type given the context and (2) play an optimal strategy against that follower type. In the remainder of this section, we will present an algorithm that meets that benchmark. In general, if the regret of an algorithm is sublinear in the time horizon $T$, i.e., $R(T)=o(T)$, we say that the algorithm is a no-regret algorithm.
@@
An important part of our analysis is that the leader be able to compute the optimal strategy for a fixed context and follower type. However, due to the tie-breaking rules of the follower, the leader might not be able to get an exact solution. Previous work [5, 24] shows that the leader can get an arbitrarily small error in the utility by optimizing over an appropriately selected subset of the simplex. For ease of notation, in the remainder of this paper, we will assume that the leader can directly optimize over the entire simplex, without that affecting her regret guarantees. For completeness, we formally state this result in Appendix A.1.
Background on Online Multiclass Learnability.
Our results build off of tools and techniques from multiclass online learning [17]. In this section, we highlight the concepts and definitions from this area that are relevant to our work. We start with the notion of shattering, which captures the following intuition: the more “complex” a set of functions is, the more complex patterns it should be able to fit.
Definition 3.3 ((Multiclass) $\pazocal{H}$-Shattered Littlestone Tree).
Let a Littlestone tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node the node-to-child edges correspond to different labels in a finite set $\mathscr{K}$. We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
The notion of multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class. [^3] Note that for binary labels, the definition of Littlestone dimension recovers the definition of the well-studied VC-dimension in the case where all nodes on the same depth of a shattered tree are the same. This definition is due to Daniely et al. [17], and corresponds to a natural generalization of the classic definition of the Littlestone dimension for binary classification [27] to the multiclass setting.
Definition 3.4 ((Multiclass) Littlestone Dimension).
@@
The finiteness of the Littlestone dimension of a hypothesis class is necessary for the class to be online learnable [27]. There are several examples of hypothesis classes, which have finite Littlestone dimension and can, therefore, be (online) learned. These include the class of conjunctions, the class of decision lists, and the class of linear thresholds with a margin (see Proposition 17 in [1]).
In Daniely et al. [17], the authors show that the multiclass Littlestone dimension of a hypothesis class characterizes the learnability of the class. In our setting, learning how to optimally predict the follower’s type given a context is an instance of multiclass prediction and is indeed sufficient for learning an optimal policy. Namely, given the context and the predicted follower type, the leader can deterministically find a strategy that maximizes her utility, by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [16]. However, it is unclear a priori whether learning to optimally predict the follower’s type is also necessary. In Theorem 3.5 we show that it is not.

### v2

@@ 3 Online Learning in Structured Stackelberg Games
In the online setting, the leader is faced with a sequence of followers. In each round $t\in[T]$, a context $\mathbf{z}_{t}\in\pazocal{Z}$ is revealed to both players. A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best-responds according to Equation˜1. We assume that the leader knows the utilities of all follower types, but the type of follower at time $t$ is not revealed to the leader until after the round is over. We make no assumptions about how the sequence of contexts is chosen, i.e., we allow them to be chosen adversarially and adaptively to the leader’s sequence of play.
3.1 Leader’s objective
Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a set of candidate functions (also called a hypothesis class) mapping contexts to follower types, which the leader uses to learn. We consider the realizable setting, where there exists a true mapping $h^{*}\in\pazocal{H}$ such that $h^{*}(\mathbf{z}_{t})=f_{t}$ for all possible $(\mathbf{z}_{t},f_{t})$ pairs. We use $h^{*}$ as a benchmark for the leader’s performance over time, which we quantify via the notion of contextual Stackelberg regret.
@@ 3.1 Leader’s objective
Given a sequence of pairs of contexts and followers $\{(\mathbf{z}_{i},f_{i})\}_{t=1}^{T}$ the leader’s regret is
$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
where $\pi_{h^{*}}(\mathbf{z})=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ and $\mathbf{x}_{t}\in\Delta(\pazocal{A})$ is the leader’s strategy at timestep $t$.
The goal of the leader is to minimize her regret, that is, to achieve utility comparable to the utility she would have gotten if she could (1) optimally predict the type of follower given the context and (2) play an optimal strategy against that type of follower. In Section˜3 we present an algorithm that is optimal with respect to this benchmark.
@@
An important part of our analysis relies on the leader being able to compute the optimal strategy for any fixed context and follower type. However, due to the tie-breaking rules of the follower, the leader might not be able to get an exact solution. Previous work  [7, 27] shows that the leader can get an arbitrarily small error in the utility by optimizing over an appropriately selected subset of the simplex. For ease of notation, in the remainder of this paper, we will assume that the leader can directly optimize over the entire simplex, without that affecting her regret guarantees. For completeness, we formally state this result in Section˜A.1.
3.2 Littlestone Dimension
In this section, we highlight the concepts and definitions from multiclass online learning [22] that are relevant to our work. We start with the notion of shattering, which captures the following intuition: the richer a set of functions is, the more complex patterns it should be able to fit.
Definition 3.3 (Multiclass $\pazocal{H}$-Shattered Littlestone Tree).
[…]
The notion of multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class. [^3] We note that for binary labels, the Littlestone dimension recovers the definition of the well-studied VC-dimension in the case where all nodes at the same depth of a shattered tree are the same. This definition is due to Daniely et al. [22] and it generalizes the classic definition of the Littlestone dimension for binary classification [30] to the multiclass setting.
Definition 3.4 (Multiclass Littlestone Dimension).
@@ 3.2 Littlestone Dimension
The finiteness of the Littlestone dimension of a hypothesis class is necessary for the class to be online learnable [30]. There are several examples of hypothesis classes that have finite Littlestone dimension and can, therefore, be (online) learned. These include the class of conjunctions, the class of decision lists, the class of linear separators with a margin (cf. Proposition 17 in [1]), and in general anything learnable in the mistake bound model.
In Daniely et al. [22], the authors show that the multiclass Littlestone dimension of a hypothesis class characterizes the learnability of the class. In our setting, learning how to optimally predict the follower’s type given a context is an instance of multiclass prediction and is indeed sufficient for learning an optimal policy. Namely, given the context and the predicted follower type, the leader can deterministically find a strategy that maximizes her utility, by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [20]. However, it is a priori unclear whether learning to optimally predict the follower’s type is also necessary. In  Theorem˜3.5 we show that the Littlestone dimension does not characterize the complexity of online learning in Stackelberg games with side information. The full proof can be found in Section˜A.2.
@@
Proof sketch.
Our proof captures the following intuition: in a structured Stackelberg game, the follower types can be completely distinct from the perspective of the leader (i.e., any one irreducible to another), and the hypothesis class can be arbitrarily hard to learn. Yet, to truly know how difficult the leader’s learning task is, one has to account for how the parameters of the game interact with the hypothesis class. Our construction uses distinct follower types and the hypothesis class indicates that there is a true threshold that splits the context domain in two parts, each of which is predictive of a different follower type. Learning that true threshold can be arbitrarily hard ($\textnormal{Ldim}(\pazocal{H})=\infty$), yet unnecessary if the two follower types happen to induce the same optimal strategy for the leader near the threshold.
This result highlights that the Littlestone dimension is completely blind to how the hypothesis class interacts with the utilities of the players, which motivates a new measure of complexity.
3.3 Stackelberg-Littlestone Dimension
In this section we provide an instance optimal learner for structured Stackelberg games. We start with the definition of node-weighted shattered trees.
[…]
Definition 3.6 (Shattered Stackelberg-Littlestone Tree).
[…]
  • (Consistency) $\pazocal{H}$ is consistent with every path in $\pazocal{T}_{d}$ i.e., for any $s\in S_{d}$, there exists $h\in\pazocal{H}$ such that for all $i\in\{\varnothing,1,\dots,d-1\}$ the label of edge $(\mathbf{z}_{s[:i]},\mathbf{z}_{s[:i+1]})$ is $h(\mathbf{z}_{s[:i]})$.
  • (Max-branch property) $\pazocal{T}_{d}$ satisfies the max-branch property, i.e., every node, $\mathbf{z}_{s}$, has children $\mathbf{z}_{si}$ for all $i\in[K]$ where $\pazocal{H}$ is consistent with path $si$.
  • (Weights) each node, $\mathbf{z}_{s}\in\mathscr{T}_{d}$ has a weight, $\rho_{s}$ defined as follows: if $\text{len}(s)=d+1$ then $\rho_{s}=0$, else
    $\rho_{s}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{sj\in S_{d}}\biggl{[}r(\mathbf{z},\mathbf{x},f^{(j)})+\rho_{sj}\biggr{]},$
    where $r(\mathbf{z},\mathbf{x},j):=u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(j)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\mathbf{x},b_{j}(\mathbf{z},\mathbf{x}))$ is the leader’s instantaneous regret, given context $\mathbf{z}$, mixed strategy $\mathbf{x}$, and follower type $j$ [^4] The infimum might not be achieved due to the followers’ tie breaking rule. We refer the reader to Section A.1 for further details.
Definition 3.7 (Stackelberg-Littlestone dimension).
The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the supremum root node weight value among all finite node-weighted trees that are shattered by hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
Note that taking the supremum is necessary since it is possible that no finite-depth tree achieves the optimal value. We begin by showing that no deterministic algorithm can achieve regret lower than the Stackelberg-Littlestone dimension.
Theorem 3.8.
For any hypothesis class, $\pazocal{H}$, any Stackelberg game $\pazocal{G}$, and any deterministic algorithm, there exists an adversarial sequence of contexts and follower types such that the learner suffers at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})-\epsilon$ regret, for any $\epsilon>0$.